OpenStudy (anonymous):

integrate 1/4e^(3x) cos(2x)

OpenStudy (anonymous):

\[\int\frac{1}{4}e^{3x}\cos(2x)~dx~?\] Integrate by parts twice.

OpenStudy (anonymous):

It's confusing me "/

OpenStudy (anonymous):

For the first round of IBP, let \[\begin{matrix} u=e^{3x}&dv=\cos(2x)~dx\\ du=3e^{3x}~dx&v=\frac{1}{2}\sin(2x) \end{matrix}\] \[\begin{align*}\frac{1}{4}\int e^{3x}\cos(2x)~dx&=\frac{1}{4}\left[\frac{1}{2}e^{3x}\sin(2x)-\frac{3}{2}\int e^{3x}\sin(2x)~dx\right]\\ &=\frac{1}{8}e^{3x}\sin(2x)-\frac{3}{8}\int e^{3x}\sin(2x)~dx \end{align*}\] For the next round, let \[\begin{matrix} u=e^{3x}&dv=\sin(2x)~dx\\ du=3e^{3x}~dx&v=-\frac{1}{2}\cos(2x) \end{matrix}\] \[\begin{align*}\frac{1}{4}\color{red}{\int e^{3x}\cos(2x)~dx}&=\frac{1}{4}\left[\frac{1}{2}e^{3x}\sin(2x)-\frac{3}{2}\int e^{3x}\sin(2x)~dx\right]\\ &=\frac{1}{8}e^{3x}\sin(2x)-\frac{3}{8}\left[-\frac{1}{2}e^{3x}\cos(2x)+\frac{3}{2}\int e^{3x}\cos(2x)~dx\right]\\ &=\frac{1}{8}e^{3x}\sin(2x)+\frac{3}{16}e^{3x}\cos(2x)-\frac{9}{16}\color{red}{\int e^{3x}\cos(2x)~dx} \end{align*}\]

OpenStudy (anonymous):

Thank you so much :D That really helped me!!!

OpenStudy (anonymous):

You're welcome!